Solve for x;
1/12 + 1/x = 1/5
Also, another problem, again solve for x
16/x-3 - 24/x+4 + 13
thank you

Answer by mahan6221(2)

1/12 + 1/x = 1/5
=> 1/x = 1/5 – 1/12
=> 1/x = (12 – 5)/60
=> 1/x = 7/60
=> x = 60/7
----------------------------
16/(x-3) – 24/(x+4) + 13
=> 16/(x-3) – 24/(x+4) + 13 = 0
=> 16(x+4)-24(x-3) + 13(x-3)(x+4) =0
=> 16(x+4)-24(x-3) + 13( x^2 + x -12) =0
=> 13(x^2) + 5x – 20 = 0
=> x = [-5 (+/-) ( 25 - 4*13*(-20))^(1/2)]/2*13
x= [-5(*/-) 32.63]/26
X1= 1.6 or X2= - 1.44